Circular No. 6574

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970111 P. J. Groot and T. Galama, University of Amsterdam (UoA); J. Telting

and R. G. M. Rutten, ING Telescopes, La Palma; and J. van Paradijs, UoA and University of Alabama in Huntsville, report: “We have obtained BVI images of the refined error-box of GRB 970111 (IAUC 6569) on Feb. 28.2 UT with integration times of 930, 405, and 695 s, respectively, and R images (405 s) on Mar. 1.1, using the 4.2-m William Herschel Telescope (prime focus) at La Palma. Comparison with the Digital Sky Survey shows no brightening of sources to V = 19. The three radio sources reported by Galama et al. (IAUC 6571) show no optical counterpart down to V = 24.0. The GRB error box contains two galaxies, at α = 15^{h}28^{m}15^{s}.2, δ = +19^{o}35′55′′ (equinox 2000.0), with V = 20.2, B − V = +1.3, V − R = +0.7, V − I = +1.4; and at α = 15^{h}28^{m}10^{s}.8, δ = +19^{o}35′35′′, with V = 20.5, B − V = +0.5, V − R = +0.4, V − I = +0.9. These galaxies are not detected in the radio observations reported by Galama et al.”

GRB 970228 Galama; R. Strom, Netherlands Foundation for Research in Astron-

omy, Dwingeloo, and UoA; van Paradijs; L. Hanlon and K. Bennett, European Space Agency/ESTEC; C. Kouveliotou, Universities Space Research Association; G. Fishman and C. Meegan, NASA Marshall Space Flight Center; J. Heise and J. in ’t Zand, SRON, Utrecht; E. Costa, M. Feroci, and L. Piro, Istituto di Astrofisica Spaziale, CNR, Frascati; F. Frontera and G. Zavattini, Universita di Ferrara; L. Nicastro and E. Palazzi, Istituto Tecnologia e Studio Radiazioni Extraterrestri, CNR, Bologna; and K. Hurley, Space Sciences Laboratory, Berkeley, report: “We made radio observations (5 GHz, 3′′.8 × 17′′ beamwidth) with the Westerbork array of the SAX/WFC error box of GRB 970228 (IAUC 6572) on Feb. 28 (19.9 hr after the onset of GRB 970228) for 1.2 hr, and on Mar. 1 and 2 for 12 hr each. The error box contains no radio sources above 0.7 mJy (4σ).”

Groot et al. (above), in ’t Zand, Heise, Costa, Feroci, Frontera, and Palazzi further report: “We have obtained V and I images of the error box of GRB 970228 (IAUC 6572) on Mar. 1.0 UT (21 hr after the BeppoSAX trigger) with integration times of 300 s each using the same telescope. Comparison with the Digital Sky Survey shows no brightening of sources to V = 19. A spiral galaxy with V = 20.5 and V − I = +1.3 is located just outside the error box at α = 5^{h}01^{m}46^{s}.0, δ = +11^{o}48′19′′ (equinox 2000.0).”

1997 March 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green